multiply and simplify
(1/3 x^4)(2/3 x^3)=
4/36x 8/27x

Answer by stanbon(75887)

multiply and simplify
(1/3 x^4)(2/3 x^3)
---
= (2/9)x^7
===================
4/36x 8/27x
---
= (1/9)x * (8/27)x
---
= (8/243)x^2
===============
Cheers,
Stan H.